Given two free vectors on a cartesian coordinate system are represented as A=3x+2y-5z and B=x-2y+z


1
Expert's answer
2021-02-23T10:03:27-0500
"\vec{A}\cdot\vec{B}=(3,2,-5)\cdot(1,-2,1)\\\vec{A}\cdot\vec{B}=-6"